### v1

# Viscosity as the product of its low-concentration value times a thermodynamic function

Abstract
The behavior of viscosity, $\eta$, as a function of concentration in dense fluids remains an unsolved problem, as is the case with other transport coefficients. Boltzmann’s theory and the Chapman-Enskog method predict the value of the viscosity at low concentrations, $\eta_{0}$. Here, the hypothesis $\eta=\phi\,\eta_{0}$ is proposed, where $\phi$ is a thermodynamic function representing the effects of interactions as concentration increases. The method proposed to verify this hypothesis involves coupling the system with a solvent represented by a Langevin thermostat, characterized by a damping time $t_{d}$. Molecular dynamics simulations show that different values of noise intensity modify $\eta$ and $\eta_{0}$, but do not affect $\phi$. This result supports the assumption that $\phi$ is a thermodynamic function, since the thermodynamic state remains unaltered by the presence of damping and noise. Simulations were conducted for particles that interact via a pseudo-hard sphere or a Lennard-Jones potential.
I Introduction
The study and understanding of transport phenomena are highly relevant to general research and to many areas of industry. Transport features of fluids can be categorized into three main types. Diffusivity relates to the displacement of particles or molecules throughout the fluid, typically composed of particles of the same type. Thermal conductivity accounts for the exchange of kinetic energy between the constituents of the fluid that arises from collisions. Finally, viscosity quantifies the friction that occurs between two adjacent layers of a fluid where the molecules move on average with different velocities. More precisely, the viscosity coefficient indicates the momentum transfer rate between layers. For dilute gases, closed expressions can be derived using the Chapman-Enskog theory Chapman and Cowling (1970), which is based on Boltzmann’s equation. An analytical description covering the entire density range has proven to be a formidable challenge. There have been several attempts to address this issue, particularly for viscosity and other transport coefficients. Enskog extended his theory to dense fluids of hard spheres Enskog (1922), where the viscosity depends on the radial distribution function $g(\sigma)$ and the second virial coefficient $b$; $\sigma$ represents the diameter of the spheres. However, this expression is only valid for small to moderate densities because one of the assumptions in the formulation is the molecular chaos approximation. Koo and Hess derived an expression for dense hard sphere fluids using the Kirkwood-Smoluchowski equation Koo and Hess (1987), which agrees with simulation results but fails at low densities. Other authors have attempted to adapt the theory of Enskog to make it suitable for real fluids Hanley et al. (1972); the result is known as Modified Enskog Theory (MET). It assumes that a real fluid generally exhibits a behavior similar to that of a hard sphere fluid, specially for small or moderate concentrations. The values of $g(\sigma)$ and $b$ are obtained from existing equations of state for real fluids. In this context, the viscosity of hard spheres, $\eta_{HS}$, is related to the Enksog viscosity, $\eta_{E}$, by the expression $\eta_{HS}=F(\rho^{*})\eta_{E}$, where $\rho^{*}$ is the reduced density and $F(\rho^{*})$ is a correction function. Additionally, the HS viscosity can be expressed as $\eta_{HS}=G(\rho^{*})\eta_{0}$, where $\eta_{0}$ is the viscosity at low concentration and $G(\rho^{*})=F(\rho^{*})\eta_{E}/\eta_{0}$. The dense gas and liquid regions have been analyzed by Dymond Dymond (1974a, b, 1985); Dymond and Assael (1996). Different expressions for the correction function have been proposed by Silva Silva et al. (2003) and by Heyes and Sigurgeirsson Heyes and Sigurgeirsson (2004). The effective Hard-Sphere Diameter method represents a real fluid as a purely repulsive hard sphere system and then considers the attractive forces as a perturbation. For instance, the Weeks-Chandler-Andersen perturbation theory (WCA) Weeks et al. (1971) splits the Lennard-Jones potential into a reference term that contains only the repulsive forces, and a perturbation term that accounts for the attractive forces that manifest in a real fluid. Silva et al. provided an approximate expression for the viscosity of a fluid with WCA interactions Silva et al. (2003)
Another approach to computing viscosity and transport coefficients in general consists of free volume theories; they provide relatively simple expressions that depend on a quantity called free volume, $V_{f}=V-V_{i}$, where $V$ is the volume and $V_{i}$ is an intrinsic molar volume. Free volume theories are generally valid only for systems with repulsive interactions, although they can be extended to include attractive interactions if the concept of activation energy is considered. Several models have been proposed, including those by Dymond-Hildebrand-Batschinski Dymond (1974a); Hildebrand (1971); Batschinski (1913), Doolittle–Cohen–Turnbull Doolittle (1951); Cohen and Turnbull (1959), Turnbull and Cohen Turnbull and Cohen (1970), and Macedo-Litovitz Macedo and Litovitz (1965).
A different method that yields good results involves excess entropy scaling laws. Rosenfeld Rosenfeld (1977a, b) derived a relation between the reduced viscosity $\eta^{*}$ and the excess entropy $S^{\text{ex}}$: $\eta^{*}=0.2\,\text{exp}(-0.8\,S^{\text{ex}}/Nk_{B})$, where $N$ is the number of particles and $k_{B}$ is the Boltzmann constant (more details on how to calculate reduced dimensionless quantities are provided below). This relation has proven to be quasi-universal, showing good agreement with simulations for various types of pair potentials, particularly in the moderate and large density regimes.
Even though we have delved into different approaches for calculating viscosity, it is worth mentioning that these methods can also be applied to obtain approximate expressions for self-diffusivity and thermal conductivity.
As we see, a significant effort has been put to provide insight into the transport process theory. So far however, finding general expressions for the transport coefficients valid for the whole density range has proven to be an arduous task. In this manuscript we aim to better understand the role that interactions play in describing viscosity for dense systems. In Ref. Marchioni et al. (2023) it was suggested that the self-diffusion coefficient $D$ may be written as $D=\varphi D_{0}$, where $D_{0}$ is the self-diffusion coefficient at low densities (which can be calculated via the kinetic theory), and $\varphi$ is a function that only depends on the thermodynamic state of the system. This implies that microscopic details such as the mass or size of the particles are not relevant for $\varphi$; instead, $\varphi$ encapsulates the macroscopic information of the interactions. Here, we seek to verify whether an equivalent relation for the shear viscosity also holds:
$\eta=\phi\,\eta_{0},$ (1)
where $\eta_{0}$ is the low-density viscosity and $\phi$ is an unknown thermodynamic function.
The procedure is analogous to that described in Marchioni et al. (2023). We assume a system of interacting particles subjected to a Langevin thermostat. That is, the particles are immersed in a solvent composed of smaller ones, which apply both a stochastic force and a damping force. The idea is that the presence of noise should not have an effect on function $\varphi$, since we assume that it is a thermodynamic function, and noise does not modify the thermodynamic state. Conversely, noise intensity is expected to affect the low concentration viscosity $\eta_{0}$, as it depends on the microscopic details of the model.
We performed molecular dynamics simulations using LAMMPS software Plimpton (1995), in which the Stoermer-Verlet time integration algorithm for constant particle number, volume, and energy (NVE ensemble) is used.
The paper is organized as follows. In Sec. II we present the motivation and justification for Eq. (1) and describe the method used to assess its validity. In Sec. III, viscosity at low concentration $\eta_{0}$ in the presence of noise is derived from the combination of Boltzmann and Langevin theories. In Sec. IV, the ratio $\eta/\eta_{0}$ for hard spheres is numerically calculated as a function of density and damping time. In Sec. V, we conduct a similar analysis for the Lennard-Jones potential at two different temperatures. Finally, conclusions are presented in Sec. VI.
II The hypotheses and the method
In Ref. Marchioni et al. (2023), it was proposed that the self-diffusion coefficient obeys a general equation of the form $D=\varphi D_{0}$; the proposal was motivated by earlier work on transition probabilities of tagged particles between two neighboring small cells Di Muro and Hoyuelos (2021). It was demonstrated Marchioni et al. (2023) that numerical simulations are consistent with the hypothesis that $\varphi$ is a thermodynamic function, independent of microscopic details. Since this is an important guide for the development of a general transport theory for dense fluids, a natural step is to verify whether s similar equation holds for other transport coefficients. Here, we propose that viscosity conforms to a similar form, as given by Eq. (1), with $\phi$ a thermodynamic function.
The method to test this hypothesis is as follows. We consider a system of $N$ particles in a volume $V$, interacting with each other via a given potential, with the temperature kept constant by a heat reservoir. Let $\mathbf{f}_{i}$ denote the force exerted on particle $i$ by the rest of the system. The particle system is immersed in a solvent, with its effect represented by a Langevin thermostat. In this setup, the total force on particle $i$ is
$\mathbf{F}_{i}=\mathbf{f}_{i}-m\mathbf{v}_{i}/t_{d}+\bm{\xi}_{i}$ (2)
where $t_{d}$ is the damping time, $m$ is the particle mass and $\mathbf{v}_{i}$ is the particle velocity; the term $-m\mathbf{v}_{i}/t_{d}$ represents friction and $\bm{\xi}_{i}$ is a stochastic force with zero mean and delta time correlated. If the temperature and the damping time are known, the noise intensity is determined by the fluctuation-dissipation theorem. The coupling to the thermostat is effectively represented by the damping time $t_{d}$.
The introduction of the thermostat modifies the transport properties; both $\eta$ and $\eta_{0}$ depend on $t_{d}$. However, the thermodynamic state remains unchanged. The equation of state is typically represented by the compressibility factor $Z=P/(k_{B}T\rho)$, where $P$ is the pressure, $k_{B}$ is Boltzmann constant, $T$ is the temperature and $\rho=N/V$ is the particle density. For given values of $T$ and $\rho$, the compressibility factor is determined by the pressure $P$. Using the virial equation (see, for example, Sec. 3.7 in Pathria and Beale (2011)), it can be shown that the pressure does not change when the thermostat is introduced (see Sec. IV in Marchioni et al. (2023) for more details); therefore, the thermodynamic state is independent of $t_{d}$. Since $\phi$ is assumed to be a thermodynamic function, it does not depend on $t_{d}$. Consequently, according to Eq. (1), $\eta/\eta_{0}$ is independent of $t_{d}$, unlike $\eta$ and $\eta_{0}$.
To verify Eq. (1), we present numerical results from molecular dynamics (MD) simulations in Sections IV and V. These simulations evaluate $\eta/\eta_{0}$ for both pseudo hard spheres and the Lennard-Jones potential across different values of $t_{d}$. The viscosity in both cases was calculated using the auto-correlation of the stress/pressure tensor integration method. The MD simulations were performed with $108000$ particles for the pseudo-hard sphere potential and $32000$ particles for the Lennard-Jones potential. We used 25 samples with an equilibration run of $10^{6}$ time steps and a data gathering run of $5\times 10^{6}$, with a time step size of $0.001$. In the next section, we will first calculate the small concentration viscosity, $\eta_{0}$, analytically.
III Viscosity at small concentration with Langevin thermostat
The small concentration viscosity can be analytically obtained using Boltzmann’s theory for the under-damped case (large $t_{d}$) and the Langevin description of Brownian motion for the over-damped case (small $t_{d}$). As discussed in the following sections, both pseudo-hard sphere and Lennard-Jones (LJ) potentials have a characteristic particle diameter, $\sigma$, and a characteristic energy, $\epsilon$. It is convenient to use reduced (dimensionless) quantities, denoted with an asterisk, which are obtained by scaling with a function of $\sigma$, $\epsilon$ or $m$. The reduced damping time is $t_{d}^{*}=t_{d}\,\sigma^{-1}\sqrt{\epsilon/m}$. Thus, a large $t_{d}$ means $t_{d}^{*}\gg 1$ and a small $t_{d}$ corresponds to $t_{d}^{*}\ll 1$.
For $t_{d}^{*}\gg 1$, the influence of the thermostat is negligible. In this case, the viscosity at low concentration can be obtained via the Boltzmann’s kinetic theory, with the Chapman-Enskog method Chapman and Cowling (1970); Hirschfelder et al. (1954),
$\eta_{B}=\frac{5}{16\,\sigma^{2}\,\Omega_{22}}\sqrt{\frac{mk_{B}T}{\pi}},$ (3)
where $\Omega_{22}$ is the so-called reduced collision integral.
The reduced temperature is given by $T^{*}=k_{B}T/\epsilon$ and the reduced viscosity is $\eta^{*}=\eta\sigma^{2}/\sqrt{m\epsilon}$. The reduced collision integral for hard spheres (HS) is equal to 1, while for the LJ potential, it has been numerically shown Fokin et al. (1999) that it can be represented by the following function of the reduced temperature:
$\Omega_{22}^{\rm LJ}(T^{*})=\exp\left(-\frac{1}{6}\log(T^{*})+\log{\frac{17}{18}}+\sum_{i=0}^{5}a_{i}{T^{*}}^{-i/2}\right),$ (4)
with $a_{0}=0.31081$, $a_{1}=-0.171211$, $a_{2}=-0.715805$, $a_{3}=2.48678$, $a_{4}=-1.78317$, $a_{5}=0.394405$. Then, we have
$\eta_{B}^{*}=\frac{5}{16\,\Omega_{22}}\sqrt{\frac{T^{*}}{\pi}}.$ (5)
On the other hand, if $t_{d}^{*}\ll 1$ the interaction with the solvent particles (the reservoir) becomes significantly more dominant than the interaction between the larger particles in the system. In this case, the dynamics of the system closely resembles that of a Brownian particle. Using Langevin theory and the Green Kubo formula, the following closed expression can be derived (see Apendix A),
$\eta_{L}=\frac{\rho\,k_{B}T\,t_{d}}{2},$ (6)
or, using reduced quantities,
$\eta_{L}^{*}=\frac{\rho^{*}\,T^{*}\,t_{d}^{*}}{2}.$ (7)
This result represents the viscosity of the system of particles, which should not be confused with the viscosity of the solvent, which is related to the damping time. In fact, the damping time and the system viscosity are inversely proportional to the viscosity of the solvent.
For intermediate values of $t_{d}^{*}$, we employ a similar approach to that presented in Ref. Marchioni et al. (2023) for the self-diffusion coefficient. Since viscosity measures the capacity for momentum transfer between different parts of a fluid, its inverse represents resistance to this transfer. Momentum transfer can be described using Newton’s law of viscosity. For example, in laminar flow in the $x$ direction, $\tau_{xy}=\eta\,\frac{\partial{u}}{\partial{y}}$, where $\tau_{xy}$ is the shear stress (the force per unit area in the $x$ direction, also interpreted as the momentum density current in the $y$ direction), and $u$ is the average velocity of the fluid in the $x$ direction. This can be analogized to Ohm’s law in an electrical circuit, that is, $I=R^{-1}\mathcal{V}$, where $I$ is the electric current, $\mathcal{V}$ is the voltage and $R$ is the electric resistance. In this analogy, we make the following correspondences: $I\rightarrow\tau$, $\mathcal{V}\rightarrow\frac{\partial{u}}{\partial{y}}$, and $R\rightarrow\eta^{-1}$. By assuming that momentum flux is analogous to the electric current in a series circuit and considering the equivalent resistance, we can derive an expression for the low concentration viscosity:
$\frac{1}{\eta_{0}^{*}}=\frac{1}{\eta_{B}^{*}}+\frac{1}{\eta_{L}^{*}},$ (8)
where $1/\eta_{B}^{*}$ and $1/\eta_{L}^{*}$ are the resistances due to collisions with other particles and due to the Langevin thermostat, respectively. Using Eqs. (5) and (7) we have,
$\eta_{0}^{*}=\frac{T^{*}\,t_{d}^{*}\,\rho^{*}}{16/5\,\rho^{*}\,t_{d}^{*}\,\Omega_{22}\sqrt{T^{*}\,\pi}+2}.$ (9)
It is straightforward to verify that $\eta_{0}^{*}\simeq\eta_{L}^{*}$ for $t_{d}\ll 1$ and that $\eta_{0}^{*}\simeq\eta_{B}^{*}$ for $t_{d}\gg 1$.
In Fig. 1 we compare Eq. (9) with numerical simulations performed at low concentration for pseudo hard-spheres ($\Omega^{\text{HS}}_{22}=1$); we observe that there is a good agreement between theory and simulations.
Figure 1: Reduced viscosity at small concentration for pseudo hard-spheres, $\eta_{0}^{*}$, against damping time, $t_{d}^{*}$, for reduced density $\rho^{*}=0.01$. Crosses are numerical results and the curve corresponds to Eq. (9).
We can see in Fig. 2 that Eq. (9) also agrees well with the results of the simulations for intermediate values of damping time and for the LJ potential, where $\Omega^{\text{LJ}}_{22}$ is given by (4).
a)Refer to caption b)Refer to caption
[…]
IV Pseudo Hard Spheres
The pairwise potential for hard-spheres (HS) is
$u_{\text{HS}}=\left\{\begin{array}[]{cc}\infty&\ \ \ r<\sigma\\[6.0pt] 0&\ \ \ r\geq\sigma\end{array}\right.,$ (10)
where $\sigma$ is the diameter of the spheres and $r$ is the distance between the centers of the spheres. It is well known that such interactions cannot be successfully implemented in Molecular Dynamics simulations, as particles that come very close to each other would experience an infinite force. However, it is possible to approximately model a hard-sphere fluid using Molecular Dynamics. Jover et al Jover et al. (2012) demonstrated that the Mie potential can effectively reproduce the properties of a hard spheres fluid. The potential is given by
$u(r)=\left\{\begin{array}[]{cc}\lambda_{r}(\frac{\lambda_{r}}{\lambda_{a}})^{\lambda_{a}}\epsilon\left[(\frac{\sigma}{r})^{\lambda_{r}}-(\frac{\sigma}{r})^{\lambda_{a}}\right]+\epsilon&\ \ \ r<\sigma\frac{\lambda_{r}}{\lambda_{a}}\\[6.0pt] 0&\ \ \ r\geq\sigma\frac{\lambda_{r}}{\lambda_{a}}\end{array}\right.,$ (11)
with $\lambda_{r}=50$ and $\lambda_{a}=49$. Note that this potential generalizes the WCA potential Weeks et al. (1971), which itself is a cut and shifted version of the well-known Lennard Jones potential. In this context, $\lambda_{r}$ and $\lambda_{a}$ are the exponents for the repulsive and attractive terms, respectively, while $\epsilon$ represents the depth of the attractive well. It has been shown that to accurately reproduce the properties of a hard-sphere fluid, the reduced temperature must be set at $T^{*}=1.5$. Since this potential does not exactly replicate a hard-core interaction, it is typically referred to as a pseudo hard-sphere potential Jover et al. (2012).
We have examined how viscosity varies with damping time at low concentration; see Eq. (9) and Fig. 1. The results indicate that, as the coupling with the reservoir decreases (i.e. as $t_{d}$ increases), the viscosity rises. At higher density, Fig. 3 illustrates the reduced viscosity as a function of reduced density for three different values of damping time (with the noise-free case corresponding to the limit of infinite $t_{d}$). As anticipated, for a given density, viscosity increases with larger $t_{d}$.
Figure 3: Viscosity for pseudo hard-spheres, $\eta^{\ast}$, versus reduced density, $\rho^{*}$, for different values of damping time $t_{d}^{*}$. The red crosses are noise-free results, blue crosses correspond to $t_{d}^{*}=10$ and green crosses to $t_{d}^{*}=100$. The circles indicate the numerical results from Pieprzyk et al. Pieprzyk et al. (2019).
Our main hypothesis is that $\eta/\eta_{0}=\phi$ is a state function. Therefore, $\phi$ should depend solely on the thermodynamic state of the system, which, as explained in Sec. II, is unaffected by the presence of noise. At low densities $\phi\simeq 1$, and it increases with concentration. However, for fixed density, changing the damping time should not modify $\phi$.
Figure 4: Viscosity divided by low-concentration viscosity for pseudo hard-spheres, $\eta/\eta_{0}$, versus reduced density, $\rho^{*}$. See Fig. 3 for a description of the symbols.
In Fig. 4 we present simulation results of $\eta/\eta_{0}=\phi$ as a function of concentration for various damping times. We find that $t_{d}$ does not influence the behavior of $\eta/\eta_{0}$. The values of $\eta/\eta_{0}$ for different damping times $t_{d}^{*}$ agree within numerical errors. For comparison, we also include numerical results from Pieprzyk et al. Pieprzyk et al. (2019) for hard spheres (without noise).
V Lennard-Jones potential
In this section we consider a pairwise interaction between particles modeled by the Lennnard-Jones potential
$u_{LJ}(r)=4\epsilon\left[\left(\frac{\sigma}{r}\right)^{12}-\left(\frac{\sigma}{r}\right)^{6}\right],$ (12)
where $\epsilon$ represents the depth of the potential well and $\sigma$ is the distance at which the potential is zero. We aim to investigate whether $\eta/\eta_{0}$ also behaves as a state function for a Lennard Jones fluid.
To test this, we first present in Fig. 5 the results of the simulations showing the reduced viscosity as a function of the concentration for three different noise intensities and two different temperatures. The results indicate that viscosity is proportional to both density and damping time.
a)Refer to caption b)Refer to caption
Figure 5: Reduced viscosity $\eta^{\ast}$ for the Lennard-Jones potential as a function of the reduced density $\rho^{\ast}$, for temperatures (a) $T=1.5$ and b) $T=4$. The red crosses are noise-free results, blue crosses correspond to $t_{d}^{*}=10$ and green crosses to $t_{d}^{*}=100$.
Next, in Fig. 6 we divide the viscosity, $\eta$, by the viscosity at low concentration, $\eta_{0}$, as given by Eq. (9); that is, $\eta/\eta_{0}=\phi$. The resulting curves for different damping times collapse, confirming the hypotheses that $\phi$ depends solely on the thermodynamic state of the system.
a)Refer to caption b)Refer to caption
Figure 6: Viscosity divided by viscosity at small concentration, $\eta/\eta_{0}$, as a function of the reduced density $\rho^{\ast}$ for the Lennard-Jones potential, with temperatures a) $T=1.5$ and b) $T=4$. See Fig. 5 for a description of the symbols.
VI Conclusions
In this work, we investigated the hypothesis that the shear viscosity of dense fluids can be expressed as $\eta=\phi\,\eta_{0}$, where $\eta_{0}$ is the low-density viscosity and $\phi$ is a thermodynamic function.
We developed a method to test this hypothesis by coupling the fluid to a Langevin thermostat characterized by a damping time $t_{d}$. This allowed us to modify $\eta$ and $\eta_{0}$ without altering the thermodynamic state of the system. An analytical expression for $\eta_{0}$ as a function of $t_{d}$ was derived by combining Boltzmann and Langevin theories. This expression showed good agreement with numerical simulations for both pseudo hard spheres and Lennard-Jones fluids. Molecular dynamics simulations were performed for pseudo hard spheres and Lennard-Jones fluids over a range of densities and damping times. The results demonstrated that while $\eta$ depends on $t_{d}$, the ratio $\eta/\eta_{0}$ is independent of $t_{d}$ within numerical uncertainty. This invariance of $\eta/\eta_{0}$ with respect to $t_{d}$ supports our hypothesis that $\phi$ is indeed a thermodynamic function, as it depends only on the thermodynamic state of the system and not on microscopic details. The behavior was consistent for both pseudo hard spheres and Lennard-Jones potentials at different temperatures, suggesting the relation may be general for simple fluids.
These results provide evidence for the proposed factorization of viscosity into a low-density term and a thermodynamic function. This formulation could simplify the development of general transport theories for dense fluids by separating the effects of dilute gas kinetics from the thermodynamic consequences of particle interactions.
Acknowledgments
This work was partially supported by Consejo Nacional de Investigaciones Científicas y Técnicas (CONICET, Argentina, PUE 22920200100016CO).
Appendix A
We consider a low-density system of Brownian particles, where interactions between particles are negligible. We aim to calculate the viscosity of this system. As mentioned in the text, it should not be confused with the viscosity of the medium in which the particles move, which is related to the friction coefficient $\gamma=1/t_{d}$, where $t_{d}$ is the damping time.
We consider a system with volume $V$, at temperature $T$, in which there are $N$ particles with mass $m$, positions $\textbf{r}_{i}=(x_{i},y_{i},z_{i})$ and velocities $\textbf{v}_{i}=(v_{xi},v_{yi},v_{zi})$, with $i=1\cdots N$. The Green-Kubo formula for the viscosity in an isotropic medium at low concentration is (see, for example, Sec. 21.9 in McQuarrie (2000))
[…]
Terms containing inter-particle forces are absent since the dilute gas limit is considered. Terms containing forces with the reservoir (friction and noise) are also neglected since it can be shown that this exchange does not contribute to the momentum flux in the system of particles.
For Brownian particles, we consider that components $x$ and $y$ of the velocity are uncorrelated, and there is also no correlation when $i\neq j$. Then,
[…]
where $\rho=N/V$ is the number density, and the isotropy property has been used.
The auto-correlation function of the velocity of Brownian particles is given by (see, for example, Sec. 5.E in Reichl (1998)),
$\langle v_{x}(0)v_{x}(t)\rangle=\frac{k_{B}T}{m}e^{-t/t_{d}}.$ (15)
Using this expression in Eq. A we obtain
$\eta_{L}=\frac{\rho\,k_{B}T\,t_{d}}{2}.$ (16)